how do you work out 1-2x=?
Answer by jai_kos(139)

Given 1-2x =
Bring the term with "x" to the other side, we get
1 = 2x
Now divide the above equation by "x", we get
1/2 =2x /2
1/2 = x
The value of "x" is 1/2